### v1

# Can we accurately read or write quantum data?

Abstract
Applications of quantum mechanics rely on the accuracy of reading and writing data. This requires accurate measurements and preparations of the quantum states. I show that accurate measurements and preparations are impossible if the total Hamiltonian is bounded from below (as thought to be in our universe). This result invites a reevaluation of the limitations of quantum control, quantum computing, and other quantum technologies dependent on the accuracy of quantum preparations and measurements, and maybe of the assumption that the Hamiltonian is bounded from below.
quantum measurement; no-go theorem; Wigner-Araki-Yanase theorem; quantum computing.
I Introduction
Quantum measurements are extremely important for all domains of the quantum, from its foundations to experiments and technological applications.
When confronted with technological limitations, for example due to errors and decoherence, we can easily get trapped in a bottom-up paradigm. The bottom-up paradigm is based on the assumption that if we solve the problems at a small scale, we can generalize the solution at a large scale. That is, it assumes that if we can build perfect quantum gates and isolate them from the environment, or if we apply quantum error correction codes, it becomes a matter of time and financial resources to apply them to build any quantum circuit and any quantum technology we can imagine on paper.
The sky seems to be the limit of quantum technologies. But up-there in the sky is Spinoza’s God, the laws of physics. Every quantum technology we have in mind has to be realized in this universe, and has to work according to the laws of physics of this universe. The physical laws may impose top-down constraints on what can be achieved. And if there are such constraints, they can’t be avoided, no matter how great researchers and engineers we have and how much time and money we invest.
We can classify the top-down problems that need to be solved in two major problems:
Problem 1 (Dynamics).
The unitary operators that we want to build as quantum gates or by combining them should be realized as restrictions of the Schrödinger unitary evolution to invariant subspaces of the total Hilbert space of our world. Since we are given only this world for this, we have to investigate what unitary operators can be realized in our world, and not merely in abstracto.
Problem 2 (Boundary conditions).
Because the technologies based on quantum information rely on reading and writing data, the accuracy of quantum measurements, either to extract information, or to prepare the quantum states, should be investigated.
Therefore, the possible limitations of quantum measurements are of particular interest. Wigner [25, 3], followed by Araki and Yanase [1], realized that additive conservation laws restrict what sharp observables can be accurately measured without being disturbed. Wigner showed that the ideal measurement of the spin along an axis is prevented by the conservation of the total angular momentum along an orthogonal axis. From this result, known as the Wigner-Araki-Yanase (WAY) theorem [14], it seems that we have to choose between inaccuracy and disturbing the system even when it is in an eigenstate of the observable. But Wigner also showed that we can avoid this problem if the measuring device is infinitely large.
Even if the same measurement cannot be both accurate and non-disturbing, for some applications it seems enough to achieve accuracy and non-disturbance in separate measurements. We can write or prepare the states using non-disturbing measurements, and read the resulting states using accurate measurements.
While Wigner concluded in [3] that accurate but not repeatable (disturbing) measurements are possible even without the requirement that the apparatus is very large, Ohira and Pearle showed that this would violate the Yanase condition [17]. The Yanase condition [26] requires the probe observable to commute with the conserved quantity, and it is necessary if we want to observe the result of the measurement.
Despite Wigner’s well known interest in symmetries, he seemed to use the existence of a conservation law as a mere signature of unitarity. But there seems to be a deeper lesson about symmetry encoded in the WAY theorem. A deep relation between symmetries, measurements, and quantum reference frames was clarified by Loveridge, Miyadera, and Busch [15]. However, the constraints on measurements extend beyond symmetries, as we can see from Ozawa’s result that measurements of non-discrete observables are not repeatable [18], and from Tukiainen’s generalization of the WAY theorem beyond conservation laws [23]. For a review of results about limitations of quantum measurements see Busch [2]. But the full scope of the WAY theorem and its generalizations is still being charted.
Ozawa warned about the possible top-down limitations imposed by the WAY theorem in [19, 20]. Lidar replied that it is already taken into account that quantum computing should be implemented on invariant subspaces of the Hilbert space [13]. More investigations of the limits imposed by conservation laws on quantum computing were reported in [12, 11].
More recently it has been shown that ideal projective measurements require infinite resource costs [9]. Without a proper preparation of the pointer state in the “ready” state, the statistics of the results will be biased. Because of the Third Law of Thermodynamics, infinite resource costs are needed for the preparation of the pointer in a suitable state, and also to ensure that the interaction between the observed system and the pointer achieves the correct correlation.
An apparently unrelated factor is the spectrum of the Hamiltonian – it is considered axiomatic to be bounded from below, i.e.  that all energy eigenvalues are larger than some fixed real value [21]. A reason often invoked is that otherwise a system will presumably decay to arbitrarily low energies [8].
Here I show that if the Hamiltonian is bounded from below, no quantum measurement and no state preparation can be accurate. The main result is proved for pure states in Section §II, and generalized to mixed states in Section §III. Section §IV concludes the article with a brief discussion of the implications of these results.
II The main result
We consider quantum measurements of an observable represented by a self-adjoint operator $\mathsf{A}$. The observed system $S$ is assumed to be initially in a state represented by the unit vector $\psi(0)$ from a Hilbert space $\mathcal{H}_{S}$. The initial, “ready” state of the measuring device $M$ is denoted by a unit vector $\varphi_{\varnothing}$ from a Hilbert space $\mathcal{H}_{M}$.
Let the measured observable $\mathsf{A}$ be, in terms of projectors $\mathsf{P}_{\lambda}$ acting on $\mathcal{H}_{S}$ and adding up to the identity,
$\mathsf{A}=\sum_{\lambda}\lambda\mathsf{P}_{\lambda}.$ (1)
Let the pointer observable of the measuring device be, in terms of projectors $\mathsf{\Pi}_{\xi}$ adding up to the identity,
$\mathsf{Z}=\sum_{\xi}\xi\mathsf{\Pi}_{\xi}.$ (2)
The observables $\mathsf{A}$ and $\mathsf{Z}$ may have continuous spectra, but for simplicity we will use the notations (1) and (2). In any case, we can decompose the continuous spectra into discrete regions and be interested only in discrete differences, since anyway in practice we discretize the results of the measurements anyway. But the proofs from this article can be applied to continuous spectra as well.
We can assume without loss of generality that the spectrum of the observable $\mathsf{Z}$ contains the spectrum of the observable $\mathsf{A}$, and an extra eigenvalue for the “ready” state $\varphi_{\varnothing}$.
The combined system consisting of the measuring device and the observed system is considered isolated for the duration of the measurement. The evolution is governed by the unitary operator $\mathbf{U}_{t}=e^{-\frac{i}{\hbar}t\mathsf{H}}$. During a time interval $[0,T]$, the two systems interact with one another. We consider interactions of the form
$\mathbf{U}_{T}\big{(}\psi(0)\otimes\varphi_{\varnothing}\big{)}=\sum_{\lambda}c_{\lambda}\psi_{\lambda}(T)\otimes\varphi_{\lambda}(T),$ (3)
where each $\varphi_{\lambda}(T)$ is an eigenstate of $\mathsf{Z}$ with eigenvalue $\lambda$. We are interested in two particular kinds of processes, measurements and preparations.
Definition 1.
An accurate measurement happens if $\mathbf{U}_{T}$ maps any separable state $\psi_{\lambda}(0)\otimes\varphi_{\varnothing}$, where $\psi_{\lambda}(0)$ is an eigenstate of $\mathsf{A}$, to another separable state $\psi_{\lambda}(T)\otimes\varphi_{\lambda}(T)$,
$\mathbf{U}_{T}\big{(}\underbrace{\psi_{\lambda}(0)}_{\text{\makebox[0.0pt]{eigenstate}}}\otimes\varphi_{\varnothing}\big{)}=\psi_{\lambda}(T)\otimes\varphi_{\lambda}(T).$ (4)
In this case, from the pointer indicating the value $\lambda$ of the pointer observable $\mathsf{Z}$, we can infer that the observed system was, immediately before the measurement, in an eigenstate of $\mathsf{A}$ with eigenvalue $\lambda$.
Definition 2.
An accurate preparation happens if, from the pointer indicating the value $\lambda$ at the end of the measurement $t=T$, we can infer that $\psi_{\lambda}(T)$ is in an eigenstate of $\mathsf{A}$ with eigenvalue $\lambda$,
$\mathbf{U}_{T}\big{(}\psi_{\lambda}(0)\otimes\varphi_{\varnothing}\big{)}=\underbrace{\psi_{\lambda}(T)}_{\text{\makebox[0.0pt]{eigenstate}}}\otimes\varphi_{\lambda}(T).$ (5)
In both cases, at time $t=T$ we retain from the superposition (3) a single outcome $\psi_{\lambda}\otimes\varphi_{\lambda}(T)$, by invoking our favorite recipe, for example the Projection Postulate or an alternative proposal based on decoherence like the many-worlds interpretation or the pilot-wave theory,
$\mathbf{U}_{T}\big{(}\psi_{\lambda}(0)\otimes\varphi_{\varnothing}\big{)}=\psi_{\lambda}(T)\otimes\varphi_{\lambda}(T)$ (6)
In both cases, we also make the following assumption, necessary for the readability of the pointer state:
Assumption 1 (Pointer stability).
There is a time interval $[T,T^{\prime}]$, $T^{\prime}>T$, so that, if the total system $S+M$ remains isolated for $t\in[T,T^{\prime}]$, the pointer states $\varphi_{\lambda}(t)$ remain eigenstates of $\mathsf{Z}$ for the eigenvalue $\lambda$, that is,
$|\varphi_{\lambda}(t)\rangle\in\Pi_{\lambda}\mathcal{H}_{M}$ (7)
for any $\lambda$, where $\Pi_{\lambda}\mathcal{H}_{M}$ is the eigenspace of $\mathsf{Z}$ corresponding to the eigenvalue $\lambda$ of $\mathsf{Z}$, cf.  equation (2).
One may imagine more general scenarios, in which the pointer indicates the state of the observed system at a different time than $t=0$ (as for reading outputs) and $t=T$ (as for reading inputs). All of these will be covered by the proof of the following result:
Theorem 1.
If the Hamiltonian is bounded from below, accurate measurements and preparations are impossible.
Proof.
We will need Lemma 1 from [10], p. 378:
Lemma 1 (Hegerfeldt & Ruijsenaars).
Let $\mathsf{H}$ be a self-adjoint operator on a Hilbert space $\mathcal{H}$, $\mathcal{H}^{\prime}$ a closed subspace of $\mathcal{H}$, and $\Psi\in\mathcal{H}$. If $\mathsf{H}$ is bounded from below and $\mathbf{U}_{t}\Psi\in\mathcal{H}^{\prime}$ for all values of $t$ in an open interval $(T,T^{\prime})$, then $\mathbf{U}_{t}\Psi\in\mathcal{H}^{\prime}$ for all $t\in\mathbb{R}$. ∎
We apply Lemma 1 to the following case (with the notations from the Lemma):
  1. 1.
    $\mathcal{H}=\mathcal{H}_{S}\otimes\mathcal{H}_{M}$.
  2. 2.
    $\Psi=\psi(0)\otimes\varphi_{\varnothing}$.
  3. 3.
    $\mathcal{H}^{\prime}=\mathcal{H}_{S}\otimes\Pi_{\lambda}\mathcal{H}_{M}$.
Indeed, from Assumption 1, $\mathbf{U}_{t}\Psi\in\mathcal{H}^{\prime}$ for all $t\in(T,T^{\prime})$, so the hypothesis of Lemma 1 is satisfied. Then, its conclusion is that $\mathbf{U}_{t}\Psi\in\mathcal{H}^{\prime}$ for all $t\in\mathbb{R}$,
$\psi(t)\otimes\varphi_{\varnothing}\in\mathcal{H}_{S}\otimes\Pi_{\lambda}\mathcal{H}_{M}\ (\forall t\in\mathbb{R}).$ (8)
In particular, for $t=0$, since $\mathbf{U}_{0}\big{(}\psi(0)\otimes\varphi_{\varnothing}\big{)}=\psi(0)\otimes\varphi_{\varnothing}$,
$\psi(0)\otimes\varphi_{\varnothing}\in\mathcal{H}_{S}\otimes\Pi_{\lambda}\mathcal{H}_{M}.$ (9)
From this,
$\varphi_{\varnothing}\in\Pi_{\lambda}\mathcal{H}_{M},$ (10)
in contradiction with the assumption that the “ready” pointer state $\varphi_{\varnothing}\in\Pi_{\varnothing}\mathcal{H}_{M}$, or, in general, that it cannot be an eigenstate of $\mathsf{Z}$ for the eigenvalue $\lambda$!
Therefore, if the Hamiltonian is bounded from below, equation (4) cannot be satisfied, so accurate measurements are impossible. Similarly, equation (5) cannot be satisfied, so accurate preparations are impossible too. ∎
Remark 1.
As defined here, measurements ensure that if the observed system is in an eigenstate of $\mathsf{A}$ with eigenvalue $\lambda$ at $t=0$, the pointer indicates the value $\lambda$, so measurements amount to reading data about the observed system. In the case of preparations, the observed system is in such an eigenstate at the end of the measurement, $t=T$, so preparations amount to writing data in the observed system’s state.
But even if the interaction between the measuring device and the observed system is such that the observed system is in the state indicated by the pointer at a different time than $t=0$ and $t=T$, this this cannot avoid the inaccuracy, because Lemma 1 implies that, if the Hamiltonian is bounded from below, equation (10) from the proof of Theorem 1 has to be true.
III Generalization to mixed states
We may hope to get around Theorem 1 by relaxing the defining conditions of measurements and preparations. For example, we can allow the observed system and the pointer to be in mixed states. Most generally, we can describe their combined state by a density operator $\rho(t)$, and not by a state vector $\Psi(t)$. Then, if the pointer state is initially in a mixed “ready” state, the initial state of the combined system satisfies the condition
$\operatorname{tr}_{S}\rho(0)=\Pi_{\varnothing}\big{(}\operatorname{tr}_{S}\rho(0)\big{)}\Pi_{\varnothing}^{\dagger},$ (11)
where $\operatorname{tr}_{S}\big{(}\rho(0)\big{)}$ is the reduced density operator obtained by partial trace over the observed system. The meaning of equation (11) is that the reduced density operator representing the pointer vanishes outside of the eigenspace representing the “ready” state of the pointer.
Condition (11) may be too relaxed, because it allows the observed system and the measuring device to be entangled from the start. If they are not entangled, $\rho(0)$ has the form $\rho(0)=\rho(0)_{S}\otimes\rho(0)_{M}$. Then already $\operatorname{tr}_{S}\rho(0)=\operatorname{tr}_{S}\big{(}\rho(0)_{S}\otimes\rho(0)_{M}\big{)}=\rho(0)_{M}$ and condition (11) becomes $\Pi_{\varnothing}\rho(0)_{M}\Pi_{\varnothing}^{\dagger}=\rho(0)_{M}$. Therefore, condition (11) is more relaxed than normally needed, but we will allow it for the sake of full generality.
After the interaction between the measuring device and the observed system, the Projection Postulate (or a decoherence-based approach) leaves us with a mixture of states containing pointer eigenstates of $\mathsf{Z}$ for only one of the eigenvalues $\lambda$. That is, it leaves us with the density operator $\big{(}\mathsf{I}_{S}\otimes\Pi_{\lambda}\big{)}\rho(T)\big{(}\mathsf{I}_{S}\otimes\Pi_{\lambda}\big{)}^{\dagger}$, where $\mathsf{I}_{S}$ is the identity operator on the Hilbert space of the observed system.
In the case of an accurate measurement, if at the beginning of the interaction the pointer state satisfied condition (11) and the observed system was a mixture of eigenstates for the eigenvalue $\lambda$, that is, it satisfied
$\operatorname{tr}_{M}\rho(0)=\mathsf{P}_{\lambda}\big{(}\operatorname{tr}_{M}\rho(0)\big{)}\mathsf{P}_{\lambda}^{\dagger}$ (12)
at the end of the interaction the total state satisfies
$\rho(T)=\big{(}\mathsf{I}_{S}\otimes\Pi_{\lambda}\big{)}\rho(T)\big{(}\mathsf{I}_{S}\otimes\Pi_{\lambda}\big{)}^{\dagger}.$ (13)
In the case of an accurate preparation, if condition (13) is satisfied (i.e.  the pointer state is a mixture of eigenstates for the eigenvalue $\lambda$ of $\mathsf{Z}$), the density operator of the observed system at the time $T$ is a mixture of eigenstates of $\mathsf{A}$ for the eigenvalue $\lambda$, so it has to vanish outside of the eigenspace of $\mathsf{A}$ with eigenvalue $\lambda$,
$\operatorname{tr}_{M}\rho(T)=\mathsf{P}_{\lambda}\big{(}\operatorname{tr}_{M}\rho(T)\big{)}\mathsf{P}_{\lambda}^{\dagger}.$ (14)
To accommodate mixed states, we relax condition (7) from Assumption 1 to
$\operatorname{tr}_{S}\rho(t)=\Pi_{\lambda}\big{(}\operatorname{tr}_{S}\rho(t)\big{)}\Pi_{\lambda}^{\dagger}$ (15)
for any $\lambda$ and $t\in[T,T^{\prime}]$.
These conditions seem to give the most relaxed, the most general characterization of accurate measurements or preparations of mixed states for sharp observables. If we try to generalize it even more by considering unsharp observables (positive operator-valued measures [7, 5]) instead of the projective observable $\mathsf{A}$, measurements or preparations will automatically be unable to accurately distinguish states, by construction.
Theorem 1 generalizes to mixed states,
Theorem 2.
If the Hamiltonian is bounded from below, accurate measurements and preparations of mixed states are impossible.
Proof.
The bounded linear operators on the Hilbert space $\mathcal{H}$ form a complex vector space $\mathcal{B}\big{(}\mathcal{H}\big{)}$, endowed with the Hilbert-Schmidt inner product, defined for any two Hermitian operators $\mathsf{B},\mathsf{C}\in\mathcal{B}\big{(}\mathcal{H}\big{)}$ by
$\big{(}\mathsf{B},\mathsf{C}\big{)}_{\textnormal{HS}}:=\operatorname{tr}\big{(}\mathsf{B}^{\dagger}\mathsf{C}\big{)}.$ (16)
This inner product is inherited by the real vector subspace $\textnormal{Herm}\big{(}\mathcal{H}\big{)}$ of bounded Hermitian operators on $\mathcal{H}$. The density operators form a convex set in $\textnormal{Herm}\big{(}\mathcal{H}\big{)}$.
The unitary evolution operators $\mathbf{U}_{t}=e^{-\frac{i}{\hbar}t\mathsf{H}}$ evolve any Hermitian operator $\mathsf{B}$ into $\mathbf{U}_{t}\mathsf{B}\mathbf{U}_{t}^{\dagger}$, preserving the Hilbert-Schmidt inner product,
$\big{(}\mathsf{B},\mathsf{C}\big{)}_{\textnormal{HS}}=\big{(}\mathbf{U}_{t}\mathsf{B}\mathbf{U}_{t}^{\dagger},\mathbf{U}_{t}\mathsf{C}\mathbf{U}_{t}^{\dagger}\big{)}_{\textnormal{HS}}.$ (17)
This applies in particular to density operators.
We define the following operators on $\mathcal{H}_{S}\otimes\mathcal{H}_{M}$,
$\widetilde{\operatorname{tr}}_{S}(\rho)$ $:=\mathsf{I}_{S}\otimes\operatorname{tr}_{S}(\rho)$ (18)
$\widetilde{\operatorname{tr}}_{M}(\rho)$ $:=\operatorname{tr}_{M}(\rho)\otimes\mathsf{I}_{M}.$
They are linear operators, because the partial traces are linear maps. Conditions (11)–(15) can be rewritten in terms of the operators $\widetilde{\operatorname{tr}}_{S}$ and $\widetilde{\operatorname{tr}}_{M}$ and the projectors
$\widetilde{\mathsf{P}}_{\lambda}$ $:=\mathsf{P}_{\lambda}\otimes\mathsf{I}_{M},$ (19)
$\widetilde{\Pi}_{\lambda}$ $:=\mathsf{I}_{S}\otimes\Pi_{\lambda},$ $\widetilde{\Pi}_{\varnothing}$ $:=\mathsf{I}_{S}\otimes\Pi_{\varnothing},$
all operating on the total Hilbert space, but also on $\textnormal{Herm}\big{(}\mathcal{H}\big{)}$, for example as in (13).
In terms of the operators (18) and (19), condition (11) becomes $\widetilde{\operatorname{tr}}_{S}\rho(0)=\widetilde{\Pi}_{\varnothing}\big{(}\widetilde{\operatorname{tr}}_{S}\rho(0)\big{)}\widetilde{\Pi}_{\varnothing}^{\dagger}$. Since the partial trace $\operatorname{tr}_{S}$ takes density operators on $\mathcal{H}_{S}\otimes\Pi_{\varnothing}\mathcal{H}$ into density operators on $\Pi_{\varnothing}\mathcal{H}$, the operator $\widetilde{\operatorname{tr}}_{S}$ takes density operators on $\mathcal{H}_{S}\otimes\Pi_{\varnothing}\mathcal{H}$ into Hermitian operators on $\mathcal{H}_{S}\otimes\Pi_{\varnothing}\mathcal{H}$. Because of this, when condition (11) is satisfied,
$\rho(0)=\widetilde{\Pi}_{\varnothing}\rho(0)\widetilde{\Pi}_{\varnothing}^{\dagger}.$ (20)
By a similar reasoning, when condition (15) is satisfied, for all $t\in(T,T^{\prime})$,
$\rho(t)=\widetilde{\Pi}_{\lambda}\rho(t)\widetilde{\Pi}_{\lambda}^{\dagger}.$ (21)
Taking into account that density operators on $\mathcal{H}$ are also vectors in the vector space $\textnormal{Herm}(\mathcal{H}\big{)}$, we apply will Lemma 1 to the following case. As a Hilbert space we use the vector space of positive semi-definite trace-class operators with finite Hilbert-Schmidt norm. Let us denote it by $\widetilde{\mathcal{H}}$. It is a real Hilbert space. Anticipating the potential objection that Lemma 1 was proved for complex, not for real Hilbert spaces, we can take the complexified $\widetilde{\mathcal{H}}\otimes\mathbb{C}$ instead of $\widetilde{\mathcal{H}}$, and extend the Hilbert-Schmidt inner product (16) to a Hermitian scalar product, obtaining a Hilbert space. Then the unitary evolution operators $\mathbf{U}_{t}$, extended by linearity to $\widetilde{\mathcal{H}}\otimes\mathbb{C}$, act unitarily on $\widetilde{\mathcal{H}}\otimes\mathbb{C}$. As the vector $\Psi$ from Lemma 1 we will take $\rho(0)$, which is a vector from the space $\textnormal{Herm}\big{(}\mathcal{H}\big{)}$. As a closed subspace $\mathcal{H}^{\prime}$ we take the subspace $\widetilde{\Pi}_{\lambda}\widetilde{\mathcal{H}}\otimes\mathbb{C}$.
Therefore, we can apply Lemma 1 in this case too, obtaining that if the Hamiltonian is bounded from below, accurate measurements and preparations are impossible even for mixed states. ∎
IV Discussion
We have seen that if the Hamiltonian is bounded from below, accurate measurements and accurate preparations are impossible. Therefore, accurate reading and writing data, being measurements and preparations, are also impossible. This doesn’t prevent us from getting asymptotically closer to perfect accuracy, just from ever reaching it. This limitation is similar to the other limitations of quantum measurements reviewed in the Introduction, but if the Hamiltonian is bounded from below it is stronger, because it prevents measurement accuracy and non-disturbance independently, and it does not need an additive conserved observable that does not commute with the observable $\mathsf{A}$.
Accuracy is important, especially if we want our technologies to be scalable and not to remain interesting experiments confined to laboratories or limited-purpose quantum computers. The resource costs to achieve even limited accuracy matter, because these resources are limited. This result affects quantum control [6], quantum computing [16], and any quantum technology dependent on the accuracy of quantum preparations and measurements, inviting us to give more careful consideration to the limits imposed by the physical laws to the technological possibilities.
Busch and Jaeger advocated for the necessity to rethink the world based on unsharp observables, to embrace the “quantum fuzziness” of the world [4]. Now we see that this becomes even more stringent.
At the same time, we may want to reconsider the assumption that our world’s Hamiltonian is bounded from below, as an alternative way to regain some of the lost certainty. This assumption leads to several results that may be problematic. First, it was shown to forbid the existence of a time operator [21], [24]. Second, it was shown to imply that relativistic wave packets confined to a bounded region of space spread instantaneously in the entire space [10]. Third, it implies that there is always a nonvanishing amplitude that the universe returns to a previous state [22]. Finally, in this article it was shown that if the Hamiltonian is bounded from below, it prevents accurate measurements and preparations. Given that all we know about the world, even its macroscopic properties, are known through observations, which are ultimately quantum, this would make everything we know about the external world to be inaccurate.